Fermi National Laboratory


Accelerator Update for July 1, 2005 - July 4, 2005

Friday July 1
The day shift began with Operations monitoring the Antiproton Source (Pbar) stack and the Recycler stash, with the Tevatron (TeV) in shot setup, and with the following experiments taking beam: MiniBooNE, NuMI.

At 8:17 AM, Operations reported that there was a problem with the Booster beam. An expert investigated.

At 8:31 AM, Operations reported a problem with the Main Injector's (MI) low-level RF system. This problem halted beam to NuMI. An operator rebooted the system and solved the problem at 8:50 AM.

Operations bypassed NuMI at 10:19 AM so that the experiment could change its target helium. The job took about two hours.

Operations established combination store 4246 at 1:14 PM with an initial luminosity of 86.63E30. Pbar resumed stacking.

Operations resumed sending beam to the Meson experiments MCenter and MTest.

At 1:31 PM, a Preaccelerator expert reported that they were going to change some memory for the H- Source. In the process of this work, they found other problems that prompted them to switch to the I- Source. Beam resumed at 2:20 PM.

At 4:37 PM, a Pbar Debuncher RF station (DRF1-3) tripped off. Operators bypassed the station, called an expert, and adjusted the RF system to compensate.

Operations put the Recycler into shot setup to stash antiprotons at 5:59 PM.

The Recycler stash ended and Pbar resumed stacking at 6:57 PM. MTest and MCenter resumed taking beam a few minutes later.

Operations reported at 8:06 PM that operators had replaced two power supplies for DRF1-3.

Operations put the Recycler into shot setup to stash antiprotons at 11:21 PM. The stash ended and Pbar resumed stacking at 11:59 PM.

Saturday July 2
The midnight shift began with Operations monitoring store 4246 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MCenter, and MTest taking beam.

At 12:08 AM, the Pbar lithium lens tripped off on a high temperature indication and wouldn't reset. Operations notified experts.

At 1:23 AM, Operations reported that a Linac RF station (KRF7) tripped off and wouldn't reset. An operator called in an expert. This shut off all downstream beam.

The Pbar experts reported that they had to replace some SCRs for the lens power supply. They completed their work by 3:53 AM.

The Linac power supply expert finished replacing a coil on KRF7. Pbar resumed stacking at 5:30 AM. MiniBooNE, NuMI, MTest, and MCenter resumed taking beam a few minutes later.

Operations terminated store 4246 at 4:05 PM and then put the TeV into shot setup ten minutes later.

At 5:34 PM, the MI power supply permit tripped off on a ground fault indication. The fault would not clear. Operators conducted a hipot, but found no ground fault. However, they located a problem with the MI-20 QF VCB. They called in experts.

The experts investigated the MI problem and reported at 9:41 PM that they found the

transformer for MI-20 QF had a ground fault. The transformer must be replaced.

Sunday July 3
The midnight shift began with Operations monitoring the stack and the Recycler stash, with MI off due to a MI-20 QF ground fault, and with MiniBooNE taking beam.

Operations prepared the TeV for access at 10:13 AM. The TeV F sector work ended at 1:41 PM.

Operations established beam to MiniBooNE at 2 PM.

Transformer Replaced An expert reported that the new MI-20 QF transformer had been placed at 3:34 PM.

At 5:41 PM, Operations reported SRM memory problems with the I-Source. Operators accessed the dome but could not reset the SRM. They found the same problem with the H- Source and determined it was a controls problem. Operators called in Linac and Controls experts.

Operations reported at 6:32 PM that the MI transformer had been placed and all connections made. On Monday a metal collar will be fabricated to encase the cables.

The Linac and controls experts brought the I- Source back on line. MiniBooNE resumed taking beam at 6:58 PM.

Operations began the TeV turn on at 7:33 PM.

Monday July 4
The midnight shift began with MI off, with Operations monitoring the stash and stack, and with MiniBooNE taking beam.

Operations put the TeV through a dry squeeze at 12:16 AM.

The Plan for Monday
The plans for today are to monitor the stack and the stash and to work on completing the MI-20 transformer hookup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/5/2005   email Fermilab